OpenStudy (trwatkins1):

What is the value of the expression –c + d when c = 12.7 and d = –4.4? A. –17.1 B. –8.6 C. 17.1 D. 8.6

OpenStudy (trwatkins1):

@pinklion23

Mehek (mehek14):

plug in the values c = 12.7 - (12.7) + (-4.4) = -12.7 + - 4.4

OpenStudy (pinklion23):

im here

OpenStudy (trwatkins1):

umm

Mehek (mehek14):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mehek14 plug in the values c = 12.7 - (12.7) + (-4.4) = -12.7 + - 4.4 \(\color{#0cbb34}{\text{End of Quote}}\) all you have to do is add 12.7 + 4.4 and put - in front of the answer

OpenStudy (pinklion23):

DELETE THAT RIGHT NOW!!!!

OpenStudy (trwatkins1):

@ganeshie8

OpenStudy (pinklion23):

i will report you!

OpenStudy (trwatkins1):

i already did

Mehek (mehek14):

did what?

OpenStudy (trwatkins1):

report

OpenStudy (pinklion23):

deez_nutz drew an inappropriate drawing

OpenStudy (pinklion23):

anyway i will help you with your question

Mehek (mehek14):

Do you still need help @trwatkins1 ?

OpenStudy (pinklion23):

oh well mehek14 did so nvm tag me in next :)

OpenStudy (trwatkins1):

ok

OpenStudy (trwatkins1):

-17.1?

Mehek (mehek14):

yes that is correct ^_^